OpenStudy (anonymous):

Evaluate -3x3 - 4x for x = -1. A. 1 B. 7 C. -1

jhonyy9 (jhonyy9):

just substitute the value of x wath is in this case -1 in place of x inside this equation ok ?

OpenStudy (anonymous):

nope i dont understand this

jhonyy9 (jhonyy9):

so will be -3x3 - 4x with x= -1 -3(-1)^3 -4(-1) = ? ok ?

OpenStudy (anonymous):

srry i dont get it

jhonyy9 (jhonyy9):

why ? there is -1 cubed = ?

OpenStudy (anonymous):

can you supply the abwser and ul bet best anwser

jhonyy9 (jhonyy9):

-3(-1)^3 -4(-1) = -3 * (-1)*(-1)*(-1) -4 * (-1) = ?

OpenStudy (anonymous):

? i dont get it

jhonyy9 (jhonyy9):

(-1) *( -1) = 1 *( -1) = -1

jhonyy9 (jhonyy9):

this wann being -1 cubed

OpenStudy (anonymous):

so it -1

jhonyy9 (jhonyy9):

just -1 cubed equal -1 but there is -3(-1)^3 -4(-1) = so than will be (-3 )*(-1) -4*(-1) = ?

OpenStudy (anonymous):

positive 1

jhonyy9 (jhonyy9):

yes than -3 * -1 = ?

OpenStudy (anonymous):

is it -1

jhonyy9 (jhonyy9):

why ? 3*1 = :

OpenStudy (anonymous):

3

jhonyy9 (jhonyy9):

yes so than you know that minus time minus make plus and that 3 time 1 equal 3 than -3 * -1 = ?

OpenStudy (anonymous):

-3

jhonyy9 (jhonyy9):

not is -3